1. Break-even analysis To be profitable, a firm must recover its costs. These costs include both its fixed and its variable cWhat affects the firms operating break-even point? Several factors affect a firms operating break-even point. Based on the
Answer #1

Breakeven units =Fixed Costs/(Sales Price-variable Costs) =12000000/(37.50-11.25) =457142.86

Number of Units*Price per unit =Fixed Costs+Breakeven Units*Variable Costs
Price per unit =(Fixed Costs+Breakeven Units*Variable Costs)/Number of Units =(12000000+175000*11.25)/175000=79.82
(Option d is correct option)

1. With increase in depreciation operating breakeven point has no change.

2. With increase in firm's tax rate operating breakeven point has no change.

3.When Variable cost decreases then operating breakeven point decreases.

Higher the operating leverage more will be its business risk.